What will be the freezing point of an aqueous solution containing 80 g of NaCl? 

Show Its Given, Required, Equation, Solution, Answer


1
Expert's answer
2022-02-14T23:47:03-0500

ΔTf=Kf×Cm


Kf(H2O)=1.86


Cm (molality) is number of moles of solute per kilogram solvent.


ΔTf=Kf×m


M (NaCl) = 58.4 g/mol


n (NaCl) = 80/58.4 = 1.37 mol


Density of water is 1 kg/l. Threfore, mass of 3.4 l of water is 3.4 kg.


Сm (NaCl) = 1.37/3.4=0.4 mol


ΔTf=1.86×0.4=0.74


Theerefore Tf of an aqueous solution containing 80.0 g of NaCl in 3.4 L of H2O will be -0.74°C.